Linear patterns of 33and 47
1
Expert's answer
2020-03-20T10:46:44-0400

To determine a linear pattern, we need to know more than just one point like (33,47). However, if the table looks like this:


"\begin{vmatrix}
x&y\\
 3 & 4 \\
 3 & 7
\end{vmatrix},"

then we can find the equation of the dependence of x upon y assuming linear pattern:

"\frac{y-y_1}{y_2-y_1}=\frac{x-x_1}{x_2-x_1},\\
\space\\
\frac{y-4}{7-4}=\frac{x-3}{3-3},\\"

therefore, the equation cannot be found. However, the "pattern" may simply look like a line with x=3 for "4\leq y\leq7":